OpenStudy (anonymous):

I want to open a bakery and I am considering locations at two different shopping centers. Use expected value to determine which location I should use. Location 1: According to my consultant, if the bakery is successful, I can anticipate an annual profit of $300,000. If it's not successful, I can anticipate an annual loss of $100,000. The probability of success at this location is 0.5. Location 2: According to my consultant, if the bakery is successful, I can anticipate an annual profit of $200,000. If it's not successful, I can anticipate an annual loss of $60,000. success is 0.75 probabilit

OpenStudy (anonymous):

@JayDS, can you help? :)

OpenStudy (anonymous):

it may be... if for expected value then for L1 300000 *0.5 =150000 as its same for L 2 200000 * 0.75 = 150000 but for loss its... L2; 60000 * 0.25 = 15000 L1 ; 100000 * 0.5= 50000 150000-15000=135000 150000-50000=100000 therefore u should chose land 2

OpenStudy (anonymous):

it can also be wrong if it is i am sorry about tht

OpenStudy (anonymous):

okay thank you:)

OpenStudy (anonymous):

check the answer plz